# Connections between Combinations Without Specified Separations and Strongly Restricted Permutations, Compositions, and Bit Strings

### v1

@@ 6 Discussion
Theorem 6, which we derived here in order to obtain the recursion relation for $B^{\{1,5\}}_{n}$, will also be used elsewhere to obtain recursion relations for $B^{\mathcal{Q}}_{n}$ and $B^{\mathcal{Q}}_{n,k}$ for four other classes of $\mathcal{Q}$; the next instance of $\mathcal{Q}$ it applies to is $\mathcal{Q}=\{1,2,5,7\}$. The metatile-generating digraph corresponding to the next case in the the $\{1,q\}$ series (Fig. 7) lacks the symmetry of the $\{1,5\}$ case and does not fall into any of the classes of digraph that we have been able to derive a recursion relation for so far. Unlike with strongly restricted permutations with the sole conditions of $\pi(i)-i\in\mathcal{D}$, where there is a general procedure for obtaining the generating function for the number of permutations [6], the literature does not appear to contain any analogous procedure for obtaining numbers of restricted combinations.
@@
There are two main features that contribute to the efficiency of the algorithm used in the program listed below. Rather than directly computing whether the difference of each pair $x,y$ of elements of subset $\mathcal{S}$ is in $\mathcal{Q}$, we perform a bitwise AND operation on bit string representations of $\mathcal{S}$ shifted $x$ places and $\mathcal{Q}$. If the AND operation gives an answer of 0 then $y-x$ does not equal any element of $\mathcal{Q}$ for all $y>x$. We also exploit the result that if $\mathcal{S}$ is an allowed subset of $\mathbb{N}_{n}$ then it is an allowed subset of $\mathbb{N}_{m}$ for any $m>n$.

### v2

@@ 6 Discussion
Theorem 6, which we derived here in order to obtain the generating function for $S^{\{1,5\}}_{n}$, will be used elsewhere to obtain generating functions for $S^{\mathcal{Q}}_{n,k}$ for various other classes of $\mathcal{Q}$. The metatile-generating digraph corresponding to the next case in the $\{1,q\}$ series (Fig. 10) does not fall into any of the classes of digraph that we have derived a generating function for so far. However, the systematic application of the digraph generating function rules (Lemmas 2, 3, and 4) can still be used to yield the generating function in this instance.
@@
There are two main features that contribute to the efficiency of the algorithm used in the program listed below. Rather than directly computing whether the difference of each pair $x,y$ of elements of subset $\mathcal{S}$ is in $\mathcal{Q}$, we perform a bitwise AND operation on bit string representations of $\mathcal{S}$ shifted $x$ places to the right and $\mathcal{Q}$. If the AND operation gives an answer of 0 then $y-x$ does not equal any element of $\mathcal{Q}$ for all $y>x$. We also exploit the result that if $\mathcal{S}$ is an allowed subset of $\mathbb{N}_{n}$ then it is an allowed subset of $\mathbb{N}_{m}$ for any $m>n$.